OpenStudy (anonymous):

What is the value of n in the equation (n – 4) – 3 = 3 – (2n + 3)?

OpenStudy (precal):

drop the parenthesis for n - 4 but distiribute -1 for 1(2n+3) move all your numbers to one side and the variable n to the other side

OpenStudy (anonymous):

wait no its 1/2 (n – 4) – 3 = 3 – (2n + 3)?

OpenStudy (precal):

ok distribute the 1/2 to (n-4)

OpenStudy (anonymous):

1/2n-2-3=6n+9

OpenStudy (precal):

right side should be 3-2n-3

OpenStudy (precal):

do you see why?

OpenStudy (anonymous):

then what would the left be

OpenStudy (precal):

your left side is correct

OpenStudy (anonymous):

so that would be 1/2n-5=-2-6 right?

OpenStudy (precal):

(1/2)n-5=-2n

OpenStudy (precal):

that is your second step

OpenStudy (anonymous):

i am soooo confused now

OpenStudy (precal):

your right side 3-(2n+3) 3-2n-3 your 3s cancel out 3-3 is zero you are left with -2n on the right side

OpenStudy (anonymous):

then wouldn't the answer be 1/2

OpenStudy (anonymous):

n=2

OpenStudy (texaschic101):

I think you understand the distribution process...you take the number outside of the parenthesis and multiply it by everything in the parenthesis. 1/2(n - 4) - 3 = 3 - (2n + 3)....keep in mind, if you just have a sign before the parenthesis, you can insert a 1. so basically your problem is : 1/2(n - 4) - 3 = 3 - 1(2n + 3) 1/2n - 2 - 3 = 3 -2n - 3.....the 3's on the right side cancel 1/2n - 5 = -2n -- subtract 1/2n from both sides -5 = -2n - 1/2n --- LCD on right side is 2 -5 = -4/2n - 1/2n -5 = -5/2n -- divide both sides by -5/2 -5/(-5/2) = n -5 * -2/5 = n 10/5 = n 2 = n check.. 1/2(2 - 4) - 3 = 3 - 1(2(2) + 3) 1/2(-2) - 3 = 3 - 1(4 + 3) -1 - 3 = 3 - 1(7) -4 = 3 - 7 -4 = -4 (correct) so n = 2 any questions at all ?